Choose all that appliable

Choose the aromatic compounds among those shown.
Answer #1
Concepts and reason

Aromatic compounds are a special type of compounds due to their structural properties like conjugation of electrons through the double bonds in a planar cyclic ring structure. Aromatic compounds are more stable. A compound can be classified as aromatic, anti-aromatic by using the Huckel rule.

Fundamentals

Specific rules for aromatic compounds are given below:

• The organic compound must be ‘cyclic’.

• The organic compound must be ‘planar’.

• The organic compound must have delocalized –electrons.

• The organic compound must have (4n+2)-electrons. Where n = 0, 1, 2, 3….

Specific rules for aromatic compounds are given below:

• The organic compound must be ‘cyclic’.

• The organic compound must be ‘planar’.

• The organic compound must have delocalized –electrons.

• The organic compound must have 4n-electrons. Where n = 0, 1, 2, 3….

The compound is given below:

CH

For the compound to be aromatic, it should satisfy the following rules.

• It should be planar;

• It should be cyclic;

• It should be delocalized –electrons; and

• It should possess the Huckel rule of 4n –electrons.

The number of –electrons = 2

On applying the Huckel rule,

4n=4n=1\begin{array}{c}\\{\rm{4n = 4}}\\\\{\rm{n = 1}}\\\end{array}

Therefore, the compound is anti-aromatic.

The compound is given below:

Oo

For the compound to be aromatic, it should satisfy the following rules:

• It should be planar;

• It should be cyclic;

• It should be delocalized –electrons; and

• It should possess the Huckel rule of 4n –electrons.

The number of –electrons = 4

On applying the Huckel rule,

4n=4n=1\begin{array}{c}\\{\rm{4n = 4}}\\\\{\rm{n = 1}}\\\end{array}

Therefore, the compound is anti-aromatic.

The compound is given below:

For the compound to be aromatic, it should satisfy the following rules:

• It should be planar

• It should be cyclic

• It should be delocalized –electrons

• It should possess the Hückel rule of 4n –electrons.

The number of –electrons = 8

On applying the Hückel rule,

4n=84n=84n=2\begin{array}{c}\\{\rm{4n = 8}}\\\\{\rm{4n = }}\frac{8}{4}\\\\{\rm{n = 2}}\\\end{array}

Therefore, the compound is anti-aromatic.

The compound is given below:

For the compound to be aromatic, it should satisfy the following rules:

• It should be planar

• It should be cyclic

• It should be delocalized –electrons

• It should possess the Hückel rule of (4n+2) –electrons.

The number of –electrons = 2

On applying the Hückel rule,

4n+2=24n=0n=0\begin{array}{c}\\{\rm{4n + 2 = 2}}\\\\{\rm{4n = 0}}\\\\{\rm{n = 0}}\\\end{array}

Therefore, the compound is aromatic.

The compound is given below:

For the compound to be aromatic, it should satisfy the following rules:

• It should be planar

• It should be cyclic

• It should be delocalized –electrons

• It should possess the Hückel rule of (4n+2) –electrons.

The number of –electrons = 6

On applying the Hückel rule,

4n+2=64n=62n=44n=1\begin{array}{c}\\{\rm{4n + 2 = 6}}\\\\{\rm{4n = 6}} - 2\\\\{\rm{n = }}\frac{4}{4}\\\\\,{\rm{n}} = 1\\\end{array}

Therefore, the compound is aromatic.

The compound is given below:

For the compound to be aromatic, it should satisfy the following rules:

• It should be planar

• It should be cyclic

• It should be delocalized –electrons

• It should possess the Hückel rule of (4n+2) –electrons.

The number of –electrons = 6

On applying the Hückel rule,

4n+2=64n=62n=1\begin{array}{c}\\{\rm{4n + 2 = 6}}\\\\{\rm{4n = 6}} - {\rm{2}}\\\\{\rm{n = 1}}\\\end{array}

Therefore, the compound is aromatic.

The compound is given below:

For the compound to be aromatic, it should satisfy the following rules:

• It should be planar

• It should be cyclic

• It should be delocalized –electrons

• It should possess the Hückel rule of (4n+2) –electrons.

The number of –electrons = 6

On applying the Hückel rule,

4n+2=64n=62n=1\begin{array}{c}\\{\rm{4n + 2 = 6}}\\\\{\rm{4n = 6}} - {\rm{2}}\\\\{\rm{n = 1}}\\\end{array}

Therefore, the compound is aromatic.

Ans:

The aromatic compounds are given below:

Aromatic
Aromatic
Aromatic
Aromatic
Anti-aromatic
Anti-aromatic
CH Anti-aromatic